### 15 Mar 2024 v3

# Internal and External Views

Abstract.
Proofs in proof assistants like Coq can be brittle, breaking easily in response to changes in the terms and types those proofs depend on. To address this, recent work introduced an algorithm and tool in Coq to automatically repair broken proofs in response to changes that correspond to type equivalences. However, many changes remained out of the scope of this algorithm and tool—especially changes in underlying behavior. We extend this proof repair algorithm so that it can express certain changes in behavior that were previously out of scope. We focus in particular on equivalences between quotient types—types equipped with a relation that describes what it means for any two elements of that type to be equal. Quotient type equivalences can be used to express interesting changes in representations of mathematical structures, as well as changes in the underlying implementations of data structures—two use cases highlighted by our case studies.
[…]
Proof Repair, Cubical Agda, Coq, Quotient Types
copyright: acmcopyrightjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYbooktitle: Woodstock ’18: ACM Symposium on Neural Gaze Detection, June 03–05, 2018, Woodstock, NYprice: 15.00isbn: 978-1-4503-XXXX-X/18/06
1. Introduction
Writing formal proofs in proof assistants like Coq, Agda, Lean, and Isabelle/HOL is a time-intensive task. Even once written, proofs may break in the face of minor changes in the datatypes, programs, and specifications they are about. User study data suggests that this process of writing and rewriting proofs is ubiquitous during proof development (Ringer et al., 2020), and that it can be challenging to deal with even for experts.
Proof repair (Ringer, 2021) aims to simplify this process by introducing algorithms and tools that fix formal proofs in response to breaking changes. In this paper, we extend proof repair to handle a new class of breaking changes in datatypes not handled by prior work. Prior work introduced a Coq plugin called Pumpkin Pi for proof repair across changes in datatypes that can be described by type equivalences (Ringer et al., 2021). Pumpkin Pi could handle only a limited class of changes corresponding to non-equivalent types by way of expressing those changes as equivalences between sigma types. The power of this was limited, however, and could not represent fundamental changes in behavior in a practical way.
In this work, we extend proof repair to support equivalences between quotient types (Section 2). Recent work in Cubical Agda showed that certain relations describing changes in behavior can be adjusted to equivalences between quotient types (Angiuli et al., 2021). However, Coq lacks quotient types entirely, so one cannot use the original Pumpkin Pi transformation (Section 3) as-is to support this class of changes. Accordingly, we explore two different views of proof repair across quotient type equivalences:
  1. (1)
    an internal view using native quotient types in Cubical Agda (Section 4), and
  2. (2)
    an external view using setoids in Coq (Section 5).
Our exploration includes two case studies in both Coq in Cubical Agda, neither of which can be handled by prior work:
  1. (1)
    We examine two non-isomorphic representations of the integers: (a) the representation used in the Cubical Agda standard library, and (b) the Grothendieck group completion of the natural numbers. We repair the addition function and proofs about it (Section 6.1).
  2. (2)
    We implement two variations of queues: (a) those backed by lists and (b) those backed by pairs of lists. These types are not equivalent, but we can construct a quotient type equivalence that describes the change and repair functions and proofs across it (Section 6.2).
Our findings are summarized in Table 1. Our code is available in supplementary material.  [^1] https://github.com/InnovativeInventor/proof-repair-quotients
Table 1. Comparison of proof repair across quotient type equivalences in Cubical Agda and Coq.
Feature Cubical Agda Coq
Quotient Types Internal External via Setoids
Repair Transformation Straightforward Extended
Correctness Some Internal External
Automation None Prototype
An Internal View in Cubical Agda
For our internal view, we adapt the transformation at the heart of the algorithm for proof repair across type equivalences to a fragment of cubical type theory, the type theory at the heart of Cubical Agda. The tradeoffs, in summary:
  1. (1)
    Quotient Types: Internal. Cubical Agda supports quotient types natively, by way of a particular higher-inductive type. Thus, we do not need to do anything special to represent quotient type equivalences; they are just equivalences between quotient types.
  2. (2)
    Repair Transformation: Straightforward. A consequence of native support for quotients is that we can adapt the transformation at the heart of proof repair across type equivalences from prior work, and use it in a straightforward way for proof repair across quotient type equivalences over our supported fragment of cubical type theory without any extensions.
  3. (3)
    Correctness: Some Internal. We can state and prove the correctness of repaired proofs internally in Cubical Agda, using cubical’s dependent path equality. Thanks to this, we are able to prove the correctness of repaired induction principles for the first time, something that was not possible in Pumpkin Pi using Coq. Still, actually composing these correctness proofs proves hard thanks to the proof-relevant notion of equality in cubical.
  4. (4)
    Automation: None. Automation in Cubical Agda is difficult due to engineering limitations and type-theoretic challenges. Our proof repair case studies in Cubical Agda are manual.
An External View in Coq
For our external view, we represent quotient types using Coq’s setoid machinery, and extend the same proof transformation to support the newly generated equality proof obligations. We also extend Pumpkin Pi with new prototype automation. The tradeoffs, in summary:
  1. (1)
    Quotient Types: External via Setoids. Coq lacks internal support for quotient types. We represent quotient types externally using Coq’s setoid machinery, and we represent quotient type equivalences as setoid equivalences.
  2. (2)
    Repair Transformation: Extended. One consequence of the lack of support for quotient types is that we must extend the proof repair algorithm to support the setoid machinery. This involves adding new cases corresponding to equality proofs and types.
  3. (3)
    Correctness: External. As in prior work, since Coq lacks univalence, we can only specify correctness by way of previously stated correctness theorems on paper inside of a univalent metatheory. We cannot state and prove them in general in Coq without introducing axioms.
  4. (4)
    Automation: Prototype. Thanks to Coq’s rich plugin infrastructure, we implement a prototype extension to the Pumpkin Pi proof repair automation in just $763$ lines of code to handle the new equality cases of the extended repair algorithm.
2. Problem Definition
The problem that we explore is an extension of the proof repair problem from Pumpkin Pi. Given proofs defined over some old type $A$, and an equivalence between $A$ and some new type $B$, Pumpkin Pi repairs proofs that refer to $A$ to instead refer to $B$. It does this by transforming those proofs in such a way that the repaired proofs no longer refer to $A$.
We expand this problem to allow for the set of pairs of equivalent types $A$ and $B$ that we can repair proofs over to include an extremely useful class of changes in types not describable in Coq—those that can be described by quotient type equivalences (Section 2.1). For simplicity, we start by considering this problem in a fragment of cubical type theory which supports quotient types directly. In such a context, we can internally and formally state what it means for any given repaired proof defined over $B$ to correctly correspond to the original proof over $A$—in terms of dependent path equality (Section 2.2). To implement this, we adapt the proof term transformation implemented in Coq for the Pumpkin Pi plugin (Section 3). We dually adapt this to support a fragment of Cubical Agda that has native support for quotient types (Section 4) and to support an extended fragment of Coq that represents quotient types using Coq’s setoid machinery (Section 5).
2.1. Scope: Quotient Type Equivalences
In this paper, we perform proof repair across quotient type equivalences. We will describe what makes these quotient type equivalences soon. But in general, a type equivalence between two types $A$ and $B$ is an isomorphism (a pair of functions that are mutual inverses) satisfying a particular coherence property (Univalent Foundations Program, 2013). It is possible to form an equivalence from any isomorphism.
Like Pumpkin Pi in Coq, we consider changes in datatypes that can be described by these type equivalences. The key difference between our work and Pumpkin Pi is that we consider a broader class of changes than type equivalences in Coq can encode. The trick is to encode changes in datatypes as equivalences between quotient types—types equipped with an equivalence relation describing what makes two elements of that type “the same.”
Quotient types are supported internally in cubical type theory by way of higher inductive types. We denote quotient types as $A/R$ for a type $A$ and a relation $R$. An element of this quotient is the equivalence class of any $a$ of type $A$. Given two elements $[a_{1}]$ and $[a_{2}]$, we have that $[a_{1}]\equiv[a_{2}]$ whenever $R\ a_{1}\ a_{2}$—even when it is not true that $a_{1}\equiv a_{2}$.
To give a familiar example of a quotient type, we can construct the natural numbers mod 2, $\mathbb{N}/2$. Two naturals are in the same equivalence class if they have the same parity. The resulting type has two equivalence classes, and so every element of the type is equal to either [ 0 ] or [ 1 ]. For another element, like [ 2 ], we have a proof of equality between [ 0 ] and [ 2 ], since both are even. This resulting type is then isomorphic to the booleans via the map [ 0 ] $\mapsto$ false, [ 1 ] $\mapsto$ true, and this can be used to construct a type equivalence.
In this paper, we further restrict ourselves to equivalences between set quotients, which have the additional constraint that all proofs of equality between elements of the quotient must themselves be equal to each other. In homotopy type theory parlance, $A/R$ must be an h-set; in broader dependent type theory parlance, uniqueness of identity proofs must provably hold for $A/R$ [^2] While homotopy type theory makes it possible to define types for which identity proofs are not unique, reasoning about equalities between those types can be challenging, so we defer reasoning about them to later work.
Quotient type equivalences let us naturally express changes in underlying implementation that break equivalence. They are useful for mathematics as well as for changes in implementation of datatypes. For example, in Section 6.2 we will show how to use quotient type equivalences to repair functions and proofs between two non-isomorphic implementations of a queue: one backed by a single list, and one backed by a pair of lists. For the latter representation, we will implement an efficient dequeueing function that operates by removing an element from the front of the second list of the pair, and an efficient enqueueing function that works by appending to the front of the first list in the pair. We will then repair proofs between these inefficient and efficient representations.
Extending the Pumpkin Pi transformation to support quotient type equivalences is what makes repair across this change possible for the first time. In Pumpkin Pi, some changes more interesting than equivalences could be expressed as equivalences between sigma types; we could try that approach for quotients, but it would come with severe disadvantages. For example, we could represent queues backed by pairs of lists as the sigma type:
$\sum_{x\>:\>\text{List }A\>\times\>\text{List }A}isCanonical\ x$
where $isCanonical\ x$ is a proposition stating that $x$ is the canonical representative of its equivalence class. This type is equivalent to our type of one list queues. However, this type has only one representative of each equivalence class present within it. With only one representative of each equivalence class present in the type, the efficient implementations of enqueue and dequeue would not be possible. Quotient types empower us to repair functions and proofs to operate over the efficient representation.
2.2. Goal: Dependent Path Equality
In cubical type theory, we can state what it means for a repaired proof to be correctly related to the original proof. This is because cubical type theory is univalent (Cohen et al., 2018; Angiuli et al., 2017)—equivalence is equivalent to equality (Univalent Foundations Program, 2013). [^3] In cubical type theory, univalence is not an axiom, but rather follows computationally from more primitive constructs. We take advantage not of univalence directly, but of these more primitive constructive constructs.
We define correctness in terms of a generalization of path equality, the primitive notion of propositional equality in cubical type theory. If two terms a and b of the same type A are path equal to each other, we can produce a term of type a $\equiv$ b.
Our two-element types from before provide an example of path equality. Using univalence, the isomorphism between $\mathbb{N}/2$ and the booleans can be turned into an equality $\mathbb{N}/2$equiv bool.
Path equality is actually an instantiation of the more powerful notion of dependent path equality, for equality between terms with different types (that is, a kind of heterogeneous equality). In this paper, we use the Cubical Agda syntax to denote dependent path equality. That is, two terms a of type A and b of type B are dependently path equal if we can produce a term of type:
  PathP p a b
where p is a path between A and B. Non-dependent path equality $\equiv$ is defined in terms of dependent path equality at the identity path. For example, let p be our proof that $\mathbb{N}/2$equiv bool. Then,
  PathP p [ 0 ] false
is inhabited, because [ 0 ] maps to false under the isomorphism we used to create p.
We can state correctness of a repaired proof relative to the corresponding original version by stating the correct PathP type—the one between the original term and the repaired term at the path between their types. The type PathP p [ 0 ] false we described above gives such an example: if we repair [ 0 ] to false across $\mathbb{N}/2\equiv$ bool, then this type states that this repair was correct.
We will describe how to define correctness more generally using PathP when we detail our internal approach in Cubical Agda in Section 4. Our external approach in Coq in Section 5 will rely solely on metatheoretical statements of correctness from Pumpkin Pi which, consistently with prior work, are not proven. One other thing to note that is also consistent with prior work is that correctness of a repaired proof relies not only on the right PathP type being inhabited, but also on the new version of the proof no longer containing references to the old version of the proof. This we do not state internally in either Cubical Agda or Coq, but it should still hold metatheoretically for repair to be correct. This is also the primary way in which proof repair in cubical type theory diverges from (dependent) transport: using transport to move proofs across equalities forces both equivalent types to eternally remain in the codebase. Repair lets us remove the old type.
3. Approach: Proof Term Transformation
To extend proof repair to support quotient type equivalences, we adapt the transformation corresponding to Pumpkin Pi’s proof repair algorithm in two different ways:
  1. (1)
    to support a fragment of cubical type theory for Cubical Agda (Section 4), and
  2. (2)
    to support an external notion of quotient types in Coq (Section 5).
The former corresponds to our internal view, while the latter corresponds to our external view.
[…]
¡t¿ ::= ¡v¿   —   ¡s¿   —   $\Pi$ (¡v¿ : ¡t¿) . ¡t¿   —   $\lambda$ (¡v¿ : ¡t¿) . ¡t¿   —   ¡t¿ ¡t¿   —   Ind (¡v¿ : ¡t¿){¡t¿,…,¡t¿}   —   Constr (¡i¿, ¡t¿)   —   Elim(¡t¿, ¡t¿){¡t¿,…,¡t¿}
[…]
    Inductive nat :=
    | O : nat
    | S : nat -> nat.
    Inductive positive :=
    | xO : positive -> positive
    | xI : positive -> positive
    | xH : positive.
    Inductive N :=
    | N0 : N
    | Npos : positive -> N.
Figure 2. The natural numbers represented in Coq, taken from the Pumpkin Pi paper (Ringer et al., 2021). The first representation is unary, and the second is binary. In positive, xH is one, xO is appending a 0 to the right side of the binary representation, and xI is appending a 1 to the right side of the binary representation. Then, N is either 0 or a positive binary number.
Pumpkin Pi calls each such decomposed equivalence a configuration, comprising pairs of the form ((DepConstr, DepElim) , ($\iota$, $\eta$)) for types on both sides of the equivalence. [^4] For the purposes of this paper, $\eta$, dealing with $\eta$-expansions of constructors applied to eliminators, will always be trivial, and thus we will ignore it. DepConstr and DepElim are, respectively, constructors and eliminators for each type. The constructors must generate the elements of the inductive type, and the eliminator must specify how to consume an element produced by the constructors. These constructors and eliminators take the shape of the original type A, even if the repaired type B has a different shape. To use an example from the Pumpkin Pi paper, for the type of unary naturals in Figure 2, we could give the following dependent constructors:
  Definition depConstrNatZero : Nat.
  Definition depConstrNatSuc : Nat -> Nat.
To repair to the binary naturals, we need to provide dependent constructors N. These dependent constructors must correspond across the equivalence. Thus, we arrive at the following constructors:
  Definition depConstrNZero : N.
  Definition depConstrNSuc : N -> N.
These dependent constructors do not share the type signatures of the type’s constructors, but rather are two user defined functions corresponding to the constructors for the unary naturals. Similarly, the dependent eliminators for both types take the same shape:
  Definition depElimNat : forall (P : nat -> Type),
    (P depConstrNatZero) -> (forall n : nat, P n -> P (depConstrNatSuc n)) ->
    forall n : nat, P n.
  Definition depElimN : forall (P : N -> Type),
    (P depConstrNZero) -> (forall n : N, P n -> P (depConstrNSuc n)) ->
    forall n : N, P n.
The fact that these two eliminators have the same shape even when the underlying types do not is exactly why we need the remaining element of the configuration: $\iota$. This gives the $\iota$-reduction rules, which specify how to reduce an application of a dependent eliminator to a dependent constructor. Over the original type, this will be definitional—the underlying proof assistant will handle it automatically. But over the repaired type, if the inductive structure has changed—as it has with binary natural numbers—the underlying proof assistant cannot handle it automatically. Instead, this reduction must be made propositional. In the binary example, $\iota$ describing how to reduce the successor case of depElimN is a propositional equality that does not hold definitionally.
flushleft $\Gamma$ $\vdash$ $t$Uparrow $t^{\prime}$
[Dep-Elim] Γ⊢a ⇑b
Γ⊢p_a ⇑p_b
Γ⊢→f_al ⇑→f_b Γ⊢DepElim(a, p_a) →f_a ⇑DepElim(b, p_b) →f_b
[Dep-Constr] Γ⊢→t_a ⇑→t_b Γ⊢DepConstr(j, A) →t_a ⇑DepConstr(j, B) →t_b
[Eta]
Γ⊢Eta(A) ⇑Eta(B)
[Iota] Γ⊢q_A ⇑q_B
Γ⊢→t_A ⇑→t_B Γ⊢Iota(j, A, q_A) →t_A ⇑Iota(j, B, q_B) →t_B
[Equivalence]
Γ⊢A ⇑B
[Constr] Γ⊢T ⇑T’
Γ⊢→t ⇑→t’ Γ⊢Constr(j, T) →t ⇑Constr(j, T’) →t’
[Ind] Γ⊢T ⇑T’
Γ⊢→C ⇑→C’ Γ⊢Ind (Ty : T) →C ⇑Ind (Ty : T’) →C’
[App] Γ⊢f ⇑f’
Γ⊢t ⇑t’ Γ⊢f t ⇑f’ t’
[Elim] Γ⊢c ⇑c’
Γ⊢Q ⇑Q’
Γ⊢→f ⇑→f’ Γ⊢Elim(c, Q) →f ⇑Elim(c’, Q’) →f’
[Lam] Γ⊢t ⇑t’
Γ⊢T ⇑T’
Γ, t : T ⊢b ⇑b’ Γ⊢λ(t : T).b ⇑λ(t’ : T’).b’
[Prod] Γ⊢t ⇑t’
Γ⊢T ⇑T’
Γ, t : T ⊢b ⇑b’ Γ⊢Π(t : T).b ⇑Π(t’ : T’).b’
[Var] v ∈Vars Γ⊢v ⇑v
Figure 3. Transformation for repair across $A\simeq B$ with configuration ((DepConstr, DepElim), (Eta, Iota)), from previous work (Ringer, 2021). Our work adapts and extends this transformation.
Once we have defined the components of the configuration, we are ready to do repair. First, the functions we wish to repair are converted to be in terms of dependent constructors, eliminators, and $\iota$-reduction rules, which Pumpkin Pi could do automatically in many cases. Then, we follow the syntactic transformation outlined in Figure 3.
We adapt this transformation for our work on proof repair across quotient type equivalences in both Cubical Agda and Coq. Note that Pumpkin Pi implements significant automation to run this transformation on existing proof terms in Coq. It does not support quotient type equivalences at all, however, since Coq does not support quotient types. Our work in Cubical Agda is manual, but supports quotient types internally; our work in Coq includes prototype automation that extends Pumpkin Pi’s automation directly, but relies on an external representation of quotient types in Coq.
4. Internal View: Proof Repair with Quotient Types in Cubical Agda
Cubical Agda supports quotient types internally. This means that, so long as we can adapt the proof term transformation to a fragment of Cubical Agda’s underlying type theory, we get proof repair across quotient type equivalences “for free,” since quotient type equivalences are just regular type equivalences (Section 4.1). Luckily, we find that adapting the algorithm to a fragment of Cubical Agda’s type system is straightforward (Section 4.2). In addition, we can leverage Cubical Agda’s dependent path equality to produce internal proofs of correctness for the individual components of proof repair (Section 4.3)—something that was not possible in Coq, and was not even done metatheoretically for any examples. Of course, one thing is notably missing from our approach in Cubical Agda: automation (Section  4.4).
4.1. Quotient Types: Internal
In Cubical Agda, set quotient types are encoded internally as higher inductive types:
  data _/_ (A : Type) (R : A normal-$\rightarrow$ A normal-$\rightarrow$ Type) : Type
    [_] : (a : A) normal-$\rightarrow$ A / R
[…]
    squash/ : (x y : A / R) normal-$\rightarrow$ (p q : x $\equiv$ y) normal-$\rightarrow$ p $\equiv$ q
For example, the type $\mathbb{N}/2$ from Section 2.1 can be captured with the following type:
  ~ : $\mathbb{N}$ $\to \mathbb{N}$ normal-$\to$ Type
  ~ n m = (n mod 2) $\equiv$ (m mod 2)
  $\mathbb{N}/2$ = $\mathbb{N}$ / $\sim$
The proof of equality between [ 0 ] and [ 2 ] is written eq/ [ 0 ] [ 2 ] refl.
Since quotient types are internal, a quotient type equivalence in Cubical Agda is simply a type equivalence that happens to be between quotient types. For the type above, we can construct an equivalence with the booleans using isoToEquiv: Iso A B normal-$\to$ A $\simeq$ B which allows us to get from any isomorphism to an equivalence. Here, we can apply isoToEquiv to the isomorphism between $\mathbb{N}/2$ and the booleans, yielding a type equivalence.
4.2. Transformation: Straightforward
[…]
We make one simplifying assumption to accommodate these set quotients. For a type T defined using set quotients, we require that the motive provided to the eliminator, (P: T normal-$\to$ Set), satisfies the condition (x: T) normal-$\to$ isSet (P x). This condition states that, for any x : T, the type P x satisfies the uniqueness of identity proofs, or, in the language of Cubical Agda, is an h-set. Because we forbid higher inductive types aside from set quotients, this condition will be satisfied for many applications in software engineering and verification.
However, this restriction does prevent the use of repair for work involving use of higher inductive types, such as the study of homotopy theory. Perhaps the most relevant limitation is that, since Set, the type of types, is not an h-set, we cannot do repair on a type family (P : T normal-$\to$ Set) such that P x is Set for some x. In the course of this work, however, we have come to believe that this simplifying assumption may not be necessary and may be removed in the future.
4.3. Correctness: Some Internal
We would like to be able to prove that our functions and proofs were repaired correctly. Intuitively, we would like for terms on the new type to behave the same as terms on the old type. However, because our old and new types are different, we cannot simply state this as a homogeneous equality. Instead, we want to know that the terms behave the same up to the equivalence we repair across.
This was done metatheoretically in a univalent type theory for the Pumpkin Pi transformation that we build off of, but it could not be done internally in general, and it was not proven for any particular example at all. In contrast, in Cubical Agda, we have access to heterogeneous path equalities in the form of PathP types. This gives us the power needed to formalize the correctness theorems about repair from the Pumpkin Pi paper in Cubical Agda, and even to prove them correct on an example type equivalence for the first time.
The shape these theorems take depends on the specific case, but the general theme is to construct a PathP between terms in the new and old type assuming the existence of such a PathP for all of its subterms. Some of these rules are generic across all types. For example, we can internally prove correctness of the Lam rule of the transformation from Figure 3 by way of functional extensionality:
[…]
    PathP ($\lambda$ i normal-$\to$ $\forall$ (t : T i) normal-$\to$ F i t) f f
[…]
Here, i, i0, and i1 are terms of the interval type, which is a primitive construct in cubical type theory from which path equalities are constructed. The rest is analogous to the transformation: f is the left function in the transformation, f is the right function, F i0 is the type of f, F i1 is the type of f, and all other subterms have the same names.
elimOK :
[…]
  $\forall$ (PA : $\mathbb{N}$ normal-$\to$ Type) (PB : Int / rInt normal-$\to$ Type) (PBSet : $\forall$ b normal-$\to$ isSet (PB b)) normal-$\to$
[…]
    PathP ($\lambda$ i normal-$\to$ Type) (PA a) (PB b)) normal-$\to$
  $\forall$ (PAO : PA zero) (PBO : PB depConstrInt/rInt0) normal-$\to$
[…]
  $\forall$ (PAS : $\forall$ a normal-$\to$ PA a normal-$\to$ PA (suc a)) (PBS : $\forall$ b normal-$\to$ PB b normal-$\to$ PB (depConstrInt/rIntS b)) normal-$\to$
[…]
      (PAS a IHa)
      (PBS b IHb)) normal-$\to$
[…]
    (Nat.elim {A = PA} PAO PAS a)
    (depElimSetInt/rInt PB PBSet PBO PBS b)
Figure 4. The theorem stating the correctness condition for the repaired dependent eliminator for a simple example type, which has been proven internally in Cubical Agda. This theorem shows that, if all the inputs to the eliminator correspond to each other across the isomorphism, then the output of the eliminator applications also corresponds across that isomorphism. Here, depConstr0OK and depConstrSOK are the correctness proofs of the repaired constructors, also proven internally.
Other rules are stated specifying the types being repaired. For example, we proved that the repaired eliminator we defined for a simple quotient equivalence was correct. Our source type was $\mathbb{N}$, and our target was Int / rInt, where Int = $\mathbb{N}\uplus\mathbb{N}$ and rInt is the relation placing inl n and inr n in the same equivalence class. The dependent constructors and eliminators correspond to the usual ones for $\mathbb{N}$. The correctness condition for a repaired eliminator for a given configuration was stated externally in the Pumpkin Pi paper, but it was not proven for any type. We adapted this theorem to Cubical Agda for our example (Figure 4) and, for the first time, proved that it held (see equivalence_int_abs.agda in the supplementary material).
One thing to note is that, even when we can internally prove individual rules of the transformation correct, it can be prohibitively challenging to actually compose those proofs to get proofs of the correctness of particular repaired proofs. Our correctness theorem is stated in terms of PathPs, but showing that two PathPs are equal requires reasoning about the equality of equality proofs between types. This means leaving the h-set fragment of cubical type theory (imposed by squash/) and entering the fragment that is proof relevant: identity proofs need not be unique, and specific proofs may be needed for goals. If our type equality proof can be written as $\lambda$ i normal-$\to$ Q (p i) where p is a path in an h-set and Q is a dependent product on elements of that h-set, we can use uniqueness of identity proofs to rewrite p to any other path with the same endpoints. But it is not always simple, or perhaps even possible, to frame the equality proof in this way.
Concretely, the way this often manifested is that we were able to compose the correctness proofs to show the correctness of repaired functions, like addition:
  addCorrect : $\forall$ (a b : $\mathbb{N}$) (a b : Int / rInt) normal-$\to$
[…]
But we could not do the same for many more interesting proofs. For example, we tried to prove that a proof of addition being commutative was repaired correctly. Trying to compose the correctness proofs of the transformation rules in the same order as the transformation rules themselves at the type level resulted in a goal of a PathP along the equality:
  $\lambda$ i normal-$\to$ addCommCorrectType
    zero depConstrInt/rInt0 depConstr0Correct
[…]
    i
But composing the correctness proofs in the natural order at the term level to attempt to fill that hole produced a PathP along the equality:
  $\lambda$ i normal-$\to$
    depConstrSCorrect i
[…]
    depConstrSCorrect i
[…]
We failed to reconcile these two equality types. This is why we say that we were able to show just some correctness proofs internally, even though in theory, it should be possible to show all of them. Fully reckoning with proof relevance so that our correctness proofs compose correctly requires additional work. For this reason, we have yet to implement internal correctness proofs in the case studies we discuss in Section 6.
4.4. Automation: None
All of our proof repair work in Cubical Agda is completely manual, even if algorithmic. This highlights a major cost of working in Cubical Agda, despite the many positive type-theoretic properties that allow us to represent quotient types natively and correctness proofs internally. Standing in the way of automation for proof repair across quotient type equivalences in Cubical Agda are both engineering and theroetical limitations. The engineering challenges include:
  1. (1)
    A lack of rich internal tooling and infrastructure. The automation in the Cubical Agda standard library operates by directly manipulating the AST of the terms, identifiers, and the type checker monad. In contrast, Coq exposes a significant amount of the internals and provides rich quality-of-life tooling for developing practical proof automation, even going as far as to provide custom debugging tooling and a custom memory allocation profiler.
  2. (2)
    No metalanguage, no side effects. In Pumpkin Pi, side effects in the metalanguage of Coq were essential for the implementation of persistent caching and other features that were necessary to build practical, performant automation. Cubical Agda does not expose a metalanguage.
  3. (3)
    Lack of documentation and examples. There are a handful of domain-specific reflective tactics currently available in the standard library, such as NatSolver and CommRingSolver. Beyond these narrow domains, there is little documentation on how to safely build more general automation in a manner that does not increase the trusted computing base, especially when it comes to interacting with the type checker monad or any of Cubical Agda’s internals.
There are also significant theoretical challenges to resolve:
  1. (1)
    Higher equalities. We do not know how to adapt automated proof repair techniques to track and discharge proof obligations unique to cubical type theory, like the interval type and boundary conditions. In particular, we do not know of a technique or decision procedure that can reason about general path equalities and higher inductive types.
  2. (2)
    Proof relevance. Cubical type theory is proof relevant. A proof relevant type theory requires one to care about the manner in which a goal was proved, a detail that poses significant challenges to building automation in Cubical Agda. Personal correspondence with Cubical Agda developers indicates hesitance to build automation infrastructure due to theoretical concerns about automation and proof relevance.
5. External View: Proof Repair with Setoids in Coq
In contrast with Cubical Agda, Coq does not support quotient types at all. So how can we support proof repair across quotient type equivalences when quotient types do not even exist? The answer is to work with setoids—an external notion of quotients in Coq (Section 5.1). We extend the Pumpkin Pi proof term transformation to work with setoids (Section 5.2). Correctness yet again becomes an external notion we can only fully express metatheoretically, as it was in prior work in Coq (Section 5.3). A notable positive is that we are able to reuse and extend Pumpkin Pi’s existing proof automation to build our prototype automation (Section 5.4).
5.1. Quotient Types: External via Setoids
In Coq, we capture quotients using setoids, which are types paired with an equivalence relation representing equality (Sozeau, 2023). For example, we can represent our $\mathbb{N}/2$ type from Section 2.1 in Coq as the setoid (nat, mod_two), where mod_two is again equivalence modulo $2$. In this case, we call nat the carrier of the setoid. Unlike in Cubical Agda, there is no equivalence class constructor, and we instead define our functions and theorems on elements of the carrier.
Coq has a Setoid type class, and any setoid defines an instance of this type class. However, our transformation will not make use of instances of Setoid. Instead, we will understand metatheoretically that any pair of a type and an equivalence relation on that type forms a setoid, and use the machinery that derives from instances of the Equivalence and Proper type classes. Notably, any type forms a setoid with the equivalence relation being equality, so all types can be consider as setoids in this way.
When comparing elements of a setoid, the equivalence relation is used in place of equality. As a consequence of this, users of setoids need to juggle multiple notions of equality, unlike with native quotient types where the same equality is used universally. To use the running example, in Cubical Agda we compare equality of elements of $\mathbb{N}/2$ using the native equality $a\equiv b$, but in Coq, when considering (nat, mod_two) as a setoid, it is generally too strong to claim that a = b. Instead, we compare equality of elements of the setoid (nat, mod_two) by using the equivalence relation mod_two a b.
Coq’s setoids do not enforce that functions defined on them respect the equivalence relation until a user needs to do rewriting under that function in a proof, unlike with native quotient types, where well-definedness is checked statically upon writing the function. For example, we can write the function:
  f (x : nat) : bool := eqb x 3
for our (nat, mod_two) setoid, and Coq allows defining this function, since it is a valid function over the carrier of the setoid. However, unlike with equality, it is not the case that f x1 = f x2 whenever mod_two x1 x2, and we could not rewrite the term f x1 to f x2 using a proof that mod_two x1 x_2. When it is the case that, for setoids (A, eqA) and (B, eqB) and a function f : A -> B, the theorem:
  forall (a1 a2 : A), (eqA a1 a2) -> eqB (f a1) (f a2)
holds, we say that f is proper. A function satisfying this property defines an instance of the Proper type class in Coq. One example of a proper function is isEven : nat -> bool, which sends even numbers to true and odd numbers to false. This is proper considering the domain as the setoid (nat, mod_two) and the codomain as the setoid (bool, =).
To do repair on Cubical Agda, we had types that were strictly equivalent after taking quotients. In Coq, we instead have a notion of a setoid equivalence. Two setoids (A, eqA) and (B, eqB) are equivalent if there is a pair of functions f : A -> B and g : B -> A satisfying the following properties:
  • f and g are proper.
  • forall (a : A), eqA (g (f a)) a
  • forall (b : B), eqB (f (g b)) b
As an example, the function isEven defined above is one half of an equivalence between (nat, mod_two) and (bool, =). It has as an inverse the function sending true to 0 and false to 1. It is easy to verify that this pair of functions satisfies the above properties.
5.2. Transformation: Extended
To adapt our transformation to setoids, we adapt our transformation to handle the changes in how equality works (Figure 5). Our transformation in Coq currently supports only the case where the source type of the transformation uses the native equality and the target type is any setoid. Because equality is an inductive type, there are three ways it manifests in terms if no axioms are used: its type, its constructors, and its eliminators.
Equality Types
For any (a1 a2 : A), there is the native type @eq A a1 a2 representing proofs of equality between a1 and a2. Every equality in the source type is translated into an equivalence relation in the target type. We require that, for each type C which lifts to a type D, the user specify an equivalence relation equiv_D and a proof that relation is an equivalence relation. In Coq, such proofs are given as instances of the Equivalence type class. If the user does not provide an equivalence relation, native equality is used instead. Then, all occurrences of @eq C are replaced with equiv_D by the transformation, as seen in the EqApp rule. Notice that we do not transform instances of @eq which are not applied to a type; this is because, until @eq is applied to some type, the transformation cannot determine which equivalence relation it should be transformed into.
Constructors
Equality has one constructor,
  @eq_refl : (A : Type) -> (a : A) -> @eq A a a
We know that @eq C will be transformed into equiv_D, and thus we must produce a term of type forall (d : D), equiv_D d d. Because the user proves that their relations are equivalence relations, each carries a proof of that term, which we will denote by reflexivity D. We can therefore replace @eq_refl C with reflexivity D, as seen in the EqReflApp rule. Again, notice that we do not repair @eq_refl unless it is applied to a type. This is for the same reason as before; until @eq_refl is applied to a type, the transformation does not know which proof of reflexivity to transform it into.
Eliminators
In Coq, equality has three eliminators, each for different sorts, but we will only focus on the Type-sorted eliminator here, @eq_rect. Equality in Coq is Leibniz equality, meaning that for any P : C -> Type, if c1 = c2 then P c1 -> P c2. Thus, the eliminators define term rewrites, with P defining where in the term rewrites take place. Specifically, for any application:
  @eq_rect (A : Type) (x : A) (P : A -> Type) (px : P x) (y : A) (H : x = y)
[…]
flushleft $\Gamma$Uparrow $\Gamma^{\prime}$
[LiftEmpty]
[LiftCons] Γ⊢x ⇑x’
Γ⊢X ⇑X’
Γ⇑Γ’ (Γ, x : X) ⇑(Γ’, x’ : X’)
[…]
[EqReflApp] Γ⊢A ⇑B Γ⊢@eq_refl (A) ⇑reflexivity(B)
[EqRewrite] Γ⊢A ⇑B
Γ⇑Γ’
Γ⊢x ⇑x’
Γ⊢P ⇑P’
Γ⊢f ⇑f’
Γ⊢y ⇑y’
Γ⊢e ⇑e’
[…]
Figure 5. The additional rules needed for repairing to a setoid. There are two mutually defined judgements. The first defines lifting of environments, and the second defines lifting of terms.
5.3. Correctness: External
[…]
Note that, even though correctness criteria like elimOK cannot in general be stated and proven internally to Coq without axioms, specific instantiations of those criteria to specific functions and proofs are sometimes provable internally. This is especially true for functions without any dependent types, for which we solely need to show that equivalent inputs map to equivalent outputs. In general, however, internalizing univalence in even an ad hoc way in Coq requires at least functional extensionality (Tabareau et al., 2021). Alternatively, one could use the homotopy type theory library for Coq, but this would rely wholly on the univalence axiom (Bauer et al., 2017).
In any case, since the proof term transformation by definition results in a proof term, that proof term can always be checked by the type checker. This means the fact that correctness is not guaranteed is not in itself too worrying. If one is solely concerned with theorems holding, and not how they are proven, then the lack of internal correctness just means that it is up to the user to check that the repaired theorem statements are the same as the original theorem statements up to the change in datatype. When one cares about the contents of functions and proofs, it is up to the user to check that those contents are the same as the original up to the change in the datatype. But the proofs produced will either hold or, in the event of a mistake in the algorithm or its implementation, simply not make it past the type checker.
5.4. Automation: Prototype
The transformation we described can be implemented in Coq using the plugin system. We have implemented a prototype of this as an extension to the Pumpkin Pi plugin. Plugins in Coq are a method of adding additional functionality to Coq. They are written in OCaml, which is the metalanguage Coq is implemented in, and can interact directly with the Coq codebase. Plugins can directly trasform and produce terms; all terms that plugins produce are checked by Coq’s type checker, and so cannot be ill typed.
Pumpkin Pi is a plugin which automatically performs proof repair across type equivalences. Pumpkin Pi has various classes of proof repair transformations across type equivaleces for which it has specialized automation. We add an additional class, termed setoid lifting, to support our extended transformation. This class mostly reuses the existing transformation, but when a source term matching the conditions set in the EqApp, EqReflApp, and EqRewrite rules is encountered, the term is translated according to those rules. In all, our extension was 763 lines of code. The implementation of these lifting rules can be found in lift.ml, liftconfig.ml, and liftrules.ml in the supplementary material.
Unlike for EqApp and EqRefl, performing the EqRewrite rule is not done as a one-for-one term substitution. Recall that we assumed the existence of an oracle which could produce proofs of rewrites. To implement this oracle, we rely on Coq’s setoid automation. Coq has a tactic, called setoid_rewrite, which attempts to perform rewriting by an equivalence relation. However, because our equivalence relations are not generally Leibniz, we must prove for each function we define that the function is proper, as defined in Section 5.1, if we wish to rewrite under applications of that function. When we prove this statement, we instantiate the Proper type class. Furthermore, the proofs that our relations are equivalences are themselves instances of the Equivalence type class. The setoid_rewrite tactic uses these type class instances to search for proofs of rewrites, and thus we can use it as our oracle.
Unlike with applications of @eq_rect, however, we cannot directly specify specific locations where we wish to perform rewrites by providing a motive P : B -> Type. To get around this, if we are trying to prove the rewrite P x -> P y, we perform a substitution P[z/y], where z is free in P, and then define P := fun z => P[z/y]. Then, for another fresh variable w, we can use the setoid_rewrite tactic to prove forall (w : B), P w x -> P w y, and recover our desired rewrite proof as (forall (w : B), P w x -> P w y) y.
Theorem depRec (C : Type)
  (posP : forall (n : nat), C)
  (negSucP : forall (n : nat), C)
  (z : GZ) :
  C.
Theorem depElimProp (P : GZ -> Prop)
  ‘(p : Proper (GZ -> Prop) (eq_GZ ==> iff) P)
  (posP : forall (n : nat), P (depConstrPos n))
  (negSucP : forall (n : nat), P (depConstrNegSuc n))
  (z : GZ) :
  P z.
Figure 6. The types of the two eliminators we use in one of our case studies. The left has non-dependently typed output, but can eliminate into Type, while the right has dependently typed output but only eliminates into Prop. The right eliminator also requires a proof that the motive is proper as a function from the setoid (Z, eq_GZ) to the setoid (Prop, iff).
To facilitate using the setoid automation, we restrict the class of objects we can repair. Specifically, whereas the original proof repair work had a single dependent eliminator, we have multiple eliminators. One eliminates into the sort Type, but is purely nondependent; in the language of Cubical Agda, this would be called a recursor. The other is dependent, but only eliminates into the sort Prop, and requires that the output motive of the eliminator be proven to be proper, considering Prop as a setoid with the if-and-only-if equivalence relation. The types of two of these eliminators for one of our case studies are in Figure 6.
The reason for this split is to facilitate automation. The setoid automation works for equivalence relations on a fixed type. As such, we cannot instantiate an instance of Equivalence for a dependently typed notion of an equivalence relation. If our eliminator into Type was dependently typed, we would then not be able to use the setoid automation to perform rewrites on applications of functions we define. For our Prop-sorted eliminator, this loss means that users cannot automatically perform rewrites on the proofs of propositions. In Coq, Prop is frequently treated as effectively proof irrelevant (despite not actually being proof irrelevant without axioms), so this loss is more acceptable.
Note that, to perform repair, we presently require that users annotate their proofs prior to running repair. This is consistent with previous work using Pumpkin Pi, where annotations were required to identify parts of the configuration to the tool, thereby decoupling the undecidable part of proof repair (configuration inference) from the decidable part (the proof term transformation itself). Even in cases where general inference is undecidable, however, Pumpkin Pi can sometimes attempt to discover these components anyway using custom heuristics for unification for particular classes of changes. We have yet to implement such inference for this class of changes. Thus, users need to explicitly use the defined dependent constructors, dependent eliminators, and iota-reduction theorems when defining terms.
Like in previous work using Pumpkin Pi, we do not directly repair terms involving pattern matching and recursion, including on equality proofs. Pumpkin Pi includes some automation to transform pattern matching and recursion to induction, which comes bundled in our extended prototype. Finally, we presently require that all instances of @eq and @eq_refl are applied to a type, and that the eliminators for @eq are fully applied.
6. Case Studies
We apply our extended proof term transformations for two proof repair case studies that use quotient type equivalences. First, we conduct repair between two representations of the integers (Section 6.1). Second, we study two common implementations of the queue data structure, and how we can repair from one to the other (Section 6.2). We do each of these case studies in both Cubical Agda (manually) and Coq (with our prototype automation when possible), demonstrating the tradeoffs of the internal and external views directly on these case studies. All of the case study examples can be found in more detail in both Cubical Agda and Coq in supplementary material.
6.1. Adding, Fast and Slow
Our first case study is mathematically motivated. We consider a change in the type representing integers. We repair addition and proofs about addition from one representation to the other. Finally, we recover the repaired proofs for a more efficient version of addition over the repaired type. Our Cubical Agda and Coq proofs for this case study can be found in grothendieck_int_equivalence.agda and grothendieck_int_equivalence.v respectively.
Types
[…]
  pos     : (n : $\mathbb{N}$) $\to \mathbb{Z}$
  negsuc : (n : $\mathbb{N}$) $\to \mathbb{Z}$
$\sim$ : ($\mathbb{N}$$\times$ $\mathbb{N}$) normal-$\to$ ($\mathbb{N}$$\times$ $\mathbb{N}$) normal-$\to$ Type
$\sim$ (x1, x2) (y1, y2) =
  x1 Nat.+ y2 $\equiv$ x2 Nat.+ y1
GZ : Type
GZ = ($\mathbb{N}$$\times$ $\mathbb{N}$) / $\sim$
Inductive Z : Set :=
| pos : nat -> Z
| negsuc : nat -> Z.
Definition GZ := nat * nat.
Definition eq_GZ (z1 z2 : GZ) :=
  match z1, z2 with
  | (a1, a2), (b1, b2) => a1 + b2 = a2 + b1
  end.
Instance eq_GZ_equiv : Equivalence eq_GZ.
Instance GZ_setoid : Setoid GZ :=
  {equiv := eq_GZ ;
  setoid_equiv := eq_GZ_equiv}.
Figure 7. The types of our integer representations in both Cubical Agda (left) and Coq (right). The Setoid instance is not explicitly needed, but is included for reader clarity.
We start with the default implementation of the integers in Cubical Agda—two copies of $\mathbb{N}$ glued together with one of them reversed, which we call $\mathbb{Z}$. We will repair functions and proofs about $\mathbb{Z}$ to use a representation of integers that may be more familiar to set theorists: Instead of two copies of $\mathbb{N}$ glued together at the ends, we consider the integers as elements of $\mathbb{N}\times\mathbb{N}/\sim$, where $(x_{1},x_{2})\sim(y_{1},y_{2})\iff x_{1}+y_{2}=x_{2}+y_{1}$. [^6] We can view this an an instance of constructing the Grothendieck group from the commutative monoid $\mathbb{N}$, and hence the integers arise as the unique group satisfying the universal property that any monoid homomorphism out of $\mathbb{N}$ can be uniquely extended to a group homomorphism out of $\mathbb{Z}$. We refer to the resulting quotient type as GZ.
The definitions of $\mathbb{Z}$ and GZ in both Cubical Agda (using quotient types) and Coq (using setoids) can be found in Figure 7. In Cubical Agda, we define the equivalence relation and then quotient $\mathbb{N}\times\mathbb{N}$ by that relation. As quotient types are internal to Cubical Agda, we are obligated to show that the relation is an equivalence relation in the construction of the quotient itself. In Coq, we still define the equivalence relation, but never formally take a quotient. Instead, we prove that the equivalence relation is an instance of the Equivalence type class in order to use Coq’s setoid rewriting automation.
In order to do repair between these two types, we need them to be isomorphic. In this case, the isomorphism is the map we expect: pos n maps to [(n , 0)], while negsuc n maps to [(0 , n + 1)]. Verifying that this map is bijective is straightforward in Cubical Agda. In Coq, because we never form equivalence classes and instead work over the carrier of the setoid, the map sends pos n to (n, 0) and negsuc n to (0 , n + 1). This map can also be seen to satisfy the definition of setoid equivalence given in Section 5.1.
Repair
To repair functions and proofs across this change, we first must decompose our isomorphism into a configuration. The full configuration for Cubical Agda can be found in Figures 15 and  16 in the appendix, while the full configuration in Coq can be found in Figures 17 and  18 in the appendix. There, we can see the dependent constructors, dependent eliminators, and $\iota$-reduction rules for both types.
The types of the components of the configuration correspond with each other, and both sides of the configuration share the same inductive structure. In Cubical Agda, the only change in types here compared to Pumpkin Pi is that every motive P : GZ normal-$\to$ Set comes with the requirement that ((x : GZ) normal-$\to$ isSet (P x)). In Coq, the configuration differs from those found in Pumpkin Pi in that there are multiple eliminators. One eliminator, which we term depRec, can only eliminate into nondependent types. The other eliminator, which we term depElimProp, can eliminate into dependent types, but only those that reside in Prop. Furthermore, to use depElimProp on GZ, we must prove that the motive P : GZ -> Prop we supply to the eliminator is a proper function, where the sort Prop is viewed as the setoid (Prop, iff). We show these eliminators in Figure 6. In addition, each of these eliminators would their own need their own set of $\iota$-reduction theorems. However, needing to $\iota$-reduce an application of depElimProp in Coq will be rare in practice, because Prop is frequently informally treated as proof irrelevant. As such, we only provide the $\iota$ rules for depRec.
Thus, with the configuration defined, we can perform repair according to the procedures discussed in Sections 4.2 and 5.2. To give one concrete example of a term being repaired, we can consider the case of repairing addition from $\mathbb{Z}$ to GZ. First, we give the definition of addition on $\mathbb{Z}$ from the Cubical Agda standard library, as well as the same algorithm implemented in Coq in Figure 8. Notice that in Cubical Agda, we can use our general eliminator, while in Coq we use our recursor.
  $\mathbb{Z}$_ : $\mathbb{Z}$ $\to \mathbb{Z}$ $\to \mathbb{Z}$
  m $\mathbb{Z}$ n =
    $\mathbb{Z}$
      ($\lambda$ $\to \mathbb{Z}$)
      ($\lambda$ p normal-$\to$ m +pos p)
      ($\lambda$ p normal-$\to$ m +negsuc p)
      n
Definition addZ (z1 z2 : Z) : Z :=
  depRecZ
    Z
    (fun (p : nat) => add_posZ z1 p)
    (fun (p : nat) => add_negsucZ z1 p)
    z2.
Figure 8. The standard definition of addition on $\mathbb{Z}$ in Cubical Agda (left) and the corresponding Coq implementation (right), explicitly annotated with the dependent eliminators.
  _+GZ_ : GZ normal-$\to$ GZ normal-$\to$ GZ
  m +GZ n =
    depElimGZ
      ($\lambda$ normal-$\to$ GZ)
      ($\lambda$ normal-$\to$ isSetGZ)
      ($\lambda$ p normal-$\to$ m +posGZ p)
      ($\lambda$ p normal-$\to$ m +negsucGZ p)
      n
Definition addGZ (z1 z2 : GZ) : GZ :=
  depRecGZ
    GZ
    (fun p : nat => add_posGZ z1 p)
    (fun p : nat => add_negsucGZ z1 p)
    z2.
Figure 9. The repaired term for addition from $\mathbb{Z}$ onto $\mathbb{Z}$, in Cubical Agda (left) and Coq (right). The repaired Cubical Agda term is manually derived, while the repaired Coq term is automatically generated.
Then, we can follow the repair algorithm to produce the repaired terms shown in Figure 9. The repair of the Cubical Agda term was performed manually. Notice that, in the Cubical Agda code, the call to $\mathbb{Z}$ in the first term is replaced with one to depElimGZ in the second term. On the other hand, repair of the Coq term was done automatically by our automation prototype by running the following command:
    Lift Z GZ in addZ as addGZ.
The transformation still behaves similarly. In the Coq code, the call to depRecZ is directly replaced with one to depRecGZ. We also see that, in the course of this repair, we had to repair two other functions: +pos and +negsuc in Cubical Agda, and their analogues add_posZ and add_negsucZ in Coq. These terms are repaired following the same algorithm, and are omitted for space.
It is worth noting that, while the functions in Coq are automatically repaired, our automation prototype does not currently attempt to generate proofs that the repaired functions are proper. Thus, the user is currently responsible for writing these proofs, though a future version of the automation could generate many of these proofs using proof search, such as Coq’s built in solve_proper tactic. Of course, for the Cubical Agda terms, this is not an issue, since it is not possible to define a function on a quotient type which is not proper.
$\mathbb{Z}$ : (z : $\mathbb{Z}$) normal-$\to$
  z $\equiv$ ($\mathbb{Z}$Pos 0) $\mathbb{Z}$ z
add0LGZ : (z : GZ) normal-$\to$
  z $\equiv$ (depConstrGZPos 0) +GZ z
add0LZ :  forall z : Z,
  z = addZ (depConstrZPos 0) z
add0LGZ : forall z : GZ,
  eq_GZ z (addGZ (depConstrGZPos 0) z)
Figure 10. A theorem whose proof we repaired, in Cubical Agda (left) and Coq (right). The Cubical Agda proof was manually repaired, while the Coq proof was automatically repaired.
We can also repair proofs. In Figure 10, we see old and new versions of a theorem whose proof we repaired (the repaired proofs are in the supplementary material). The theorem states that 00 is a left identity for addition. Again, we can see elements of one side of the configuration swapped for the other. However, we also see that in Coq, equality has been replaced with an equivalence relation on the type, reflecting the fact that we are making a claim about setoid equality.
As before, the repair on Cubical Agda terms was manual, and was automated for Coq. However, the current version of Pumpkin Pi does not support configurations with multiple eliminators, and proving add0LZ requires using both depRecZ and depElimPropZ. We circumvent this by first defining constants referring to all subterms which contain depRecZ, and then separately repairing those constants. Then, we rewrite the proof of the theorem using these constants and depElimPropZ. We then reconfigure Pumpkin Pi to use depElimPropZ (by calling Pumpkin Pi’s configuration command a second time), and then repair the rewritten theorem after reconfiguring. This is not a fundamental limitation of our approach, and is inherited from Pumpkin Pi; future versions can remove the need for this workaround by supporting multiple eliminators in a single configuration.
In addition to the mentioned terms, we also repair the successor and predecessor functions on integers, functions which add $n$ or subtract $n+1$ from an integer (which were used above to define general addition), and the proof of a theorem saying that 00 is a right identity for addition. In Coq, all these functions and proofs could be repaired automatically, while repair was conducted manually in Cubical Agda. All of these functions and proofs, as well as their repaired versions, can be found in the supplementary material.
Fast and Slow
Proof repair allows us to repair functions defined on our type, which are then used in proofs. However, the operation of the repaired functions reflects the inductive structure of the old type, rather than that of the new type. Accordingly, the repaired functions are often inefficient.
Here, this manifests as follows: The most general proof that we repair from $\mathbb{Z}$ to GZ is the eliminator for $\mathbb{Z}$, which we use to define both addition and the proof about it. This gives us a repaired eliminator for GZ, which occurs inside of the repaired addition function and proof over GZ. But this repaired eliminator is slow, and so are our repaired functions and proofs that use it. It is not the eliminator that one would naturally define for GZ. Rather, it internally computes a canonical representative of the equivalence class of that element, either of the form $(n,0)$ for $n\geq 0$ or $(0,n)$ for $n\geq 1$. Because these canonical elements operate like pos n and negsuc n from the original type, we can compute on them in the same way as we did the original type—in other words, it follows the inductive structure of $\mathbb{Z}$. This means that, for any computation, we must compute canonical elements of equivalence classes, which is wasteful.
Instead, we would much rather use the more natural function $\lambda(n_{1},n_{2})(m_{1},m_{2}).(n_{1}+m_{1},n_{2}+m_{2})$. We are able to define this by way of the term in Figure 11.
addHelpFunc : ($\mathbb{N}\times\mathbb{N}$) normal-$\to$ ($\mathbb{N}\times\mathbb{N}$) normal-$\to$ ($\mathbb{N}\times\mathbb{N}$)
[…]
addResp : (a a b b : $\mathbb{N}\times\mathbb{N}$) normal-$\to$
  $\sim$ a a normal-$\to$
  $\sim$ b b normal-$\to$
  $\sim$ (addHelpFunc a b) (addHelpFunc a b’)
-- proof of addResp omitted for space
addGZ : GZ normal-$\to$ GZ normal-$\to$ GZ
addGZ =
  setQuotBinOp
    isReflR
    isReflR
    addHelpFunc
    addResp
Definition fastAddGZ (a b : GZ) :=
  match b with
  | (b1, b2) =>
    match a with
    | (a1, a2) => (a1 + b1, a2 + b2)
    end
  end.
Figure 11. Our fast addition function on the repaired integers, in Cubical Agda (left) and Coq (right). In addGZ’, setQuotBinOp is a specialized eliminator that allows us to easily define a binary function on set quotients. The definition of fast addition in Coq is shorter because we do not need to prove respectfulness when defining it. The use of pattern matching is acceptable because this function is neither to be repaired nor a product of repair.
The proofs we repaired are for the repaired addition function, as opposed to this more natural and efficient addition function. Thankfully, in Cubical Agda, we are able to use functional extensionality to show that the slow repaired function and the fast addition function are equal, since they have the same output for any inputs:
  addEqual : addGZ $\equiv$ _+GZ_
  addEqual = funExt ($\lambda$ x normal-$\to$ funExt ($\lambda$ y normal-$\to$ addEqualOnInputs x y))
Then, to obtain proofs for our fast addition function, we merely need to substitute the fast function for the repaired function in the proofs—something we can do easily thanks to transport. We can then repair our example proof:
  add’0LGZ : (z: GZ) normal-$\to$ z $\equiv$ addGZ (depConstrGZPos 0) z
  add’0LGZ = subst
    ($\lambda$ y normal-$\to$ (z : GZ) normal-$\to$ z $\equiv$ y (depConstrGZPos 0) z)
    (sym addEqual)
    add0LGZ
We can move between slow and fast functions like this in Coq as well, but (consistently with prior work) we must take a more ad hoc approach, since Coq’s type theory has neither functional extensionality nor transport. Instead, we can define our fastAddGZ function, and prove the theorem:
  Theorem addEqualFastAdd : forall (a b : GZ), eq_GZ (addGZ a b) (fastAddGZ a b).
Then, as long as we can rewrite all applications of addGZ in a theorem statement, we can obtain the corresponding theorem about fastAddGZ. For example, we can translate add0LGZ into the theorem:
  Theorem fastAdd0LGZ : forall (z : GZ), eq_GZ z (fastAddGZ (depConstrGZPos 0) z).
using only one rewrite by addEqualFastAdd. If applications of addGZ were inside opaque terms, however, we may not be able to view all the applications of addGZ, and thus could not rewrite like this. Higher order functions are one potential source of this trouble; for instance, if the definition of List.map is opaque, and we pass addGZ as the function to map over a list, we would not be able to access the application sites of addGZ and thus could not rewrite by addEqualFastAdd. If we knew that fastAddGZ = addGZ, we could perform the rewrite anyway, which is one advantage Cubical Agda’s approach holds over Coq.
6.2. Variations on a Theme of Queues
Next, we repair functions and proofs across a change in implementation of a queue data structure. This is motivated by an example from Angiuli et al. (2021), which showed that quotient types can be use to adjust certain relations more general than equivalences into equivalences for use with transport in Cubical Agda. That class of changes was cited in the Pumpkin Pi paper as an example that could not be expressed naturally in Coq with the original framework. In Cubical Agda, we can express this use case naturally, highlighting the benefits of working in a type theory with quotient types. In Coq, with our extensions to Pumpkin Pi, we can express this using setoids. Our Cubical Agda and Coq proofs for this case study can be found in two_list_queue_equivalence.agda and two_list_queue_equivalence.v, respectively.
OLQ = List A
~ : (List A $\times$ List A) normal-$\to$
      (List A $\times$ List A) normal-$\to$
     Type
~ (l1, l2) (l3, l4) =
  l1 ++ (rev l2) = l3 ++ (rev l4)
TLQ = (List A $\times$ List A) / $\sim$
Definition OLQ := list A.
Definition TLQ := list A * list A.
Definition insOrder (q : TLQ) :=
  match q with
  | (l1, l2) => l1 ++ rev l2
  end.
Definition eq_queue (q1 q2 : TLQ) :=
  insOrder q1 = insOrder q2.
Instance eq_queue_equiv : Equivalence eq_queue.
Instance TLQ_setoid : Setoid TLQ  :=
  {equiv := eq_queue ;
  setoid_equiv := eq_queue_equiv}.
Figure 12. One list queues (OLQ) and two list queues (TLQ) in Cubical Agda (left) and Coq (right). The Setoid instance is not explicitly needed, but is included for reader clarity.
Types
Our first implementation represents queues using a single list, given by the type OLQ in Figure 12. The intention behind this representation is that elements enter the queue at the front of the list and are removed from the queue at the back of the list. The simplicity of this representation of queues comes with a cost: our dequeue operation runs in linear time.
Our second representation is more complicated, but resolves the runtime issues. Instead of one list, we use a two list representation of queues as List A $\times$ List A. Here, elements enter the queue by being added to the front of the first list, and are removed from the queue by being removed from the front of the second list. If the second list is empty, the first list is reversed onto the second list.
To do repair between these types, we need them to be isomorphic, but the types List A and (List A $\times$ List A) are not naturally isomorphic. Based on our description, we would say that a two list queue (l1 , l2) corresponds to the queue l1 ++ (rev l2), where ++ is the list append operator. However, this is not an injective map; multiple two list queues correspond to a single one list queue. In Cubical Agda, we resolve this by taking a quotient of our type of two list queues. We define the equivalence relation (l1 , l2) $\sim$ (l3 , l4) $\iff$ l1 ++ (rev l2) = l3 ++ (rev l4) and quotient (List A $\times$ List A) by it, obtaining the type TLQ of two list queues seen in Figure 12. The resulting two types have an isomorphism by the function [(l1 , l2)] $\mapsto$ (l1 ++ (rev l2)), which is well defined on our quotient type and has as an inverse the function l $\mapsto$ [(l , [])]. In Coq, we do not have quotient types, but instead work in the setoid on (list A * list A) with the previously mentioned equivalence relation.
Repair
We wish to conduct repair from our one list queue type to our two list queue type. We first write the configuration for our isomorphism. The Cubical Agda configuration can be found in Figures 19 and  20, while the Coq configuration can be found in Figures 21 and  22, both of which are in the appendix. We have the standard functions enqueue and dequeue defined for one list queues, and we wish to repair them to our newly defined two list queue type. We also have theorems about these functions whose proofs we want to repair. The main theorem states that enqueue and dequeue are related in the way we expect (Figure 13).
  returnOrEnq : A normal-$\to$ Maybe (OLQ $\times$ A) normal-$\to$ OLQ $\times$ A
  returnOrEnq a =
    Cubical.Data.Maybe.rec
    (depConstrEmpty , a)
    ($\lambda$ p normal-$\to$
[…]
  dequeueEnqueue : (a : A) (q : OLQ) normal-$\to$
    dequeue (enqueue a q) $\equiv$
    just (returnOrEnq a (dequeue q))
Definition returnOrEnq (a : A) (m : option (OLQ * A)) : (OLQ * A) :=
  @option_rect
    (queue * A)
    (fun _ => prod queue A)
    (fun (p : (queue * A)) =>
      (enqueue a (fst p), snd p))
    (depConstrEmpty, a)
    m.
Theorem dequeueEnqueue (a : A) (q : OLQ) :
  (dequeue (enqueue a q)) =
  (Some (returnOrEnq a (dequeue q))).
Figure 13. Main theorem statement that relates dequeue and enqueue in Cubical Agda (left) and Coq (right), given over one list queues.
We then repair our functions and proofs. To do this, first we annotate our functions and proofs to use the dependent constructors, dependent eliminators, and $\iota$-reduction rules. (Pumpkin Pi partially automates this process for some classes of changes (though the general problem is undecidable), but as previously mentioned, we do not yet support this for the setoid class of changes in our prototype extension.) Then, we follow the transformation outlined in Sections 4.2 and 5.2 to repair these terms. In Cubical Agda, the terms are repaired manually. In Coq, the terms are repaired automatically by our Pumpkin Pi extension. For example, dequeueOLQ was repaired by running the command
    Lift OLQ TLQ in dequeueOLQ as dequeueTLQ.
As explained previously, however, the user still needs to prove that the lifted functions are proper. Also, in this case we need to manage multiple equivalence relations, since the return type of dequeueTLQ is option (TLQ * A) rather than simply TLQ. The user must define equivalence relations for all types they use which need a relation other than equality and provide them to Pumpkin Pi.
As an example, we repair the dequeueOLQ function on queues in Cubical Agda and Coq. The definitions for this functions, along with their repaired versions, can be found in the appendix. (Figure 23 and Figure 24). Like before, the dependent constructors, dependent eliminators, and $\iota$-reduction steps for our one list queues are replaced with those for two list queues in our dequeue function. In addition to this, we also repaired the functions enqueueOLQ, dequeueHelpOLQ, and the proofs dequeueEmptyOLQ and dequeueEnqueueTLQ to two list queues, thus providing the standard queue API and its specification, in both Cubical Agda and Coq. All of these terms were repaired manually in Cubical Agda, and were repaired automatically in Coq.
Fast and Slow
While we have obtained a dequeue operation on two list queues, the repaired implementation is inefficient. Our repaired dequeue, as a consequence of the implementation of our dependent constructors and eliminators, always returns the representative of the equivalence class with all elements on the first list. This is undesirable, since we have to do extra computation to move elements from the second list to the first list. The Coq version is similarly inefficient. We would much prefer dequeue be implemented in the way originally described, simply taking the first element off the second list in the pair. We can implement this dequeue function in Cubical Agda and Coq (Figure 14).
fastDequeue : TLQ normal-$\to$ Maybe (TLQ $\times$ A)
fastDequeue = SetQuotients.rec isSetDeqReturnType func wellDefined where
func : TLQ normal-$\to$ Maybe (TLQ $\times$ A)
func ([] , []) = nothing
func ((a :: l1) , []) =
  just (_/_.[ [] ,
    safe-tail (rev (a :: l1)) ] ,
    safe-head a (rev (a :: l1)))
func ([] , (a :: l2)) =
  just (_/_.[ [] , l2 ], a)
func ((b :: l1) , (a :: l2)) =
  just (_/_.[ (b :: l1) , l2 ] , a)
  -- proof of well-definedness omitted
Definition fastDequeueTLQ (q : TLQ) :=
  let (l1, l2) := q in
  match l1, l2 with
  | [] , [] => None
  | h1 :: t1 , [] =>
    Some (([] , tl (rev l1)), hd h1 (rev l1))
  | _ , h2 :: t2 =>
    Some ((l1, t2), h2)
  end.
Figure 14. Fast dequeue function for two list queues in Cubical Agda (left) and Coq (right).
Now, we want to know that our repaired dequeue is equal to fastDequeue. That way, any theorems we prove about dequeue on one list queues can easily be applied to fastDequeue by substituting over the equality. In Cubical Agda, with functional extensionality, we obtain a proof of the equality:
  deqIsFastDeq : dequeue $\equiv$ fastDequeue
With this theorem, we know that anywhere our repaired dequeue appears, we can instead use our more efficient fastDequeue. In Coq, as before, we cannot show that the functions are equal, but we can show they are equal pointwise:
  Theorem dequeueEqualsFastDequeue : forall (q : TLQ),
    eq_deq_ret (dequeueTLQ q) (fastDequeueTLQ q).
Here, eq_deq_ret is the user-provided equivalence relation on option  (TLQ * A). Then, we can rewrite applications of dequeueTLQ to applications of fastDequeueTLQ, and translate theorems about dequeueTLQ to fastDequeueTLQ that way.
7. Related Work
Proof Repair. This work is based on the proof repair work by Ringer et al. (2021). This work examines how repair can be adapted to support quotient type equivalences, a class of changes previously not supported. We adapt and extend their proof term transformation as well as the Pumpkin Pi Coq plugin that implements automation for that transformation. We also explore the applicability to a fragment of Cubical Agda by way of a manual repair process. Pumpkin Pi has some more mature automation for other classes of changes, like automatic search for configurations, that we cannot yet support in our prototype extension for quotient type equivalences.
Proof repair was first introduced in parallel by Ringer et al. (2018) and Robert (2018), with strong influence from the field of program repair (Monperrus, 2017). Sisyphus (Gopinathan et al., 2023) is a recent proof repair tool that, like our work, can handle changes in behavior (using a mix of dynamic and static techniques). However, Sisyphus repairs proofs of imperative OCaml programs verified in Coq using an embedded separation logic, whereas our work repairs proofs that are written in Cubical Agda and Coq directly.
Proof Reuse. Proof repair is an instance of proof reuse, which seeks to use existing proofs in new goals. Other work in proof reuse includes CoqEAL (Cohen et al., 2013) which uses refinement relations to verify properties of efficient functions using proofs on functions that are easy to reason about. CoqEAL can handle relations more general than equivalences, but does not include support for porting proofs across those changes. In Isabelle/HOL, the Transfer package (Huffman and Kunčar, 2013) uses automation to transfer proofs between types. Both approaches require the source and target type to remain in the codebase, unlike proof repair. A complementary approach is to design proofs to be more reusable or more robust to changes from the start (Woos et al., 2016; Chlipala, 2013; Delaware et al., 2011). More work on proof reuse can be found in the QED at Large (Ringer et al., 2019) survey of proof engineering.
Quotients and Equivalences. Our work uses quotient types to expand the scope of proof repair across type equivalences. Quotient types exist in other proofs assistants besides Cubical Agda, like Isabelle/HOL (Wenzel et al., 2004; Isabelle Development Team, 2024), as well as Lean (Avigad et al., 2017) by way of axioms. Bortin and Lüth (2010) use quotient types to construct theories in Isabelle, such as multisets and finite sets as quotients of lists. Coq does not have quotient types, but it does have setoids (Sozeau, 2023), which do not explicitly form equivalence classes like quotients do. Setoid type theory uses a setoid model to justify the axioms needed to represent quotient types (Altenkirch et al., 2019). We draw on proper quotient types for our internal work for Cubical Agda, and we draw on setoids for our external work for Coq.
Our idea for extending proof repair using quotient type equivalences to begin with comes from Angiuli et al. (2021), which shows that certain relations more general than equivalences can be represented this way. The first example present in that paper is the queue example which we have also studied in our work. Because that work uses transport, it requires the user to keep both versions of the type in their codebase. We avoid that problem, but also have to reason more closely about the inductive structure of our types. In doing so, we extend proof repair to support a new class of changes described as missing from the original Pumpkin Pi work in Ringer et al. (2021).
Univalent Foundations. This project was conducted partially in Cubical Agda, and partially assuming a univalent metatheory. Cubical Agda is an implementation of cubical type theory (Vezzosi et al., 2019). Cubical type theory (Cohen et al., 2018; Coquand et al., 2018; Angiuli et al., 2017) was developed to give a constructive account of the univalence axiom. When working in Cubical Agda, we are able to state and prove internal correctness of parts of our repair transformation and have a computational interpretation of functional extensionality. Cubical type theory itself is a derivative of Voevodsky’s homotopy type theory (Univalent Foundations Program, 2013), which presents the univalence axiom non-constructively. Homotopy type theory has additionally been implemented in Coq as the HoTT library (Bauer et al., 2017).
Work has been done to approximate univalence in Coq. Tabareau et al. (2018) defines univalent parametricity, which allows the transport of a restricted class of functions and theorems. Univalent parametricity implements an ad hoc form of transport that only sometimes requires the axiom of functional extensionality, and in many cases is axiom-free. It also includes a form of type-directed search to transport terms by way of type classes, something that proof repair tools like Pumpkin Pi and our extension still lack. Subsequent work introduces a white-box transformation (Tabareau et al., 2021) similar to the repair transformation from Pumpkin Pi, which Ringer (2021) describes as developed in parallel with mutual influence. None of these support quotient type equivalences like our work does, though it is possible that by leaning further on the axiom of functional extensionality, one could use these tools with quotient type equivalences as well.
8. Conclusions & Future Work
We extended proof repair across type equivalences to support changes represented by quotient type equivalences, adding more expressive power. This power stemmed from the insight from prior work in cubical type theory that certain relations more general than equivalences can be viewed as equivalences between quotient types. Realizing this insight concretely in a proof repair context made it possible to support changes never before supported by a proof repair tool.
The key challenge we overcame was supporting quotient types in a proof repair algorithm initially built for a type theory that does not even have quotient types to begin with. Our internal approach addressed this by adapting the entire algorithm to a fragment of cubical type theory, then manually porting functions and proofs in Cubical Agda using that algorithm. Our external approach addressed this by representing quotient types using Coq’s setoids, extending the existing algorithm and automation to dispatch the newly generated equality proof obligations, and running the newly extended automation to port functions and proofs in Coq. The former shined in its internal support of quotient types, its internal correctness proofs, and its corresponding simplicity; the latter shined in its automation and corresponding relative ease of use.
Going forward, we hope to realize all of these advantages in the same tool—something we could do easily if only we knew of a univalent language with native quotient types that also has a strong metalanguage for building automation. We hope to continue to improve our automation and its usability well beyond a prototype. In an internal context, we hope to better compose the internal correctness proofs in a way that could be automated in a proof-relevant type theory. In an external context, we hope to look at other kinds of types and relations that can be expressed even when the type theory lacks them, as quotient types can be by way of setoids. We hope all of this will open the door to supporting proof repair across more and more sophisticated classes of changes.
Acknowledgements
We thank Amélia Liao, Reed Mullanix, Tom Jack, and the Cubical Agda Discord server for their help in using Cubical Agda. We thank also Carlo Angiuli for their early thoughts on this project. This research was developed with funding from the Defense Advanced Research Projects Agency. The views, opinions, and/or findings expressed are those of the author(s) and should not be interpreted as representing the official views or policies of the Department of Defense or the U.S. Government.
Appendix A Complete configuration for integer case study
In order to perform repair between our two representations of integers, we must first define a configuration to perform the repair across. In Figures 15 and  16, we see the types of the elements of the configuration defined in Cubical Agda. The types on the left make up the part of the configuration relating to our inductive definition of integers, and the types on the right make up the part of the configuration relating to our quotient representation of the integers. We align the types to highlight that the types on both sides have the same shape, except for the hypothesis ((x : GZ) normal-$\to$ isSet (P x)) in depElimGZ. Note that the Cubical Agda configuration only has one eliminator.
In Figures 17 and  18, we see the types of the elements of the configuration for repair in Coq. Again, the types on the left correspond with our inductively defined integer representation, and the types on the right correspond with our setoid representation of the integers. The types are again aligned to highlight the similarity in the shapes of both sides, with the only difference being that depElimPropGZ takes a proof of:
    ‘(p : Proper (GZ -> Prop) (eq_GZ ==> iff) P).
as an argument. In this configuration, there are two eliminators. The first eliminates into Type, but does not support dependently typed output, while the second only eliminates into Prop but does support dependently typed output.
$\mathbb{Z}$Pos : $\mathbb{N}$ $\to \mathbb{Z}$
$\mathbb{Z}$Pos = pos n
$\mathbb{Z}$NegSuc : $\mathbb{N}$ $\to \mathbb{Z}$
$\mathbb{Z}$NegSuc = negsuc n
$\mathbb{Z}$ : (P : $\mathbb{Z}$ normal-$\to$ Type) normal-$\to$
  ((n : $\mathbb{N}$) normal-$\to$ P ($\mathbb{Z}$Pos n)) normal-$\to$
  ((n : $\mathbb{N}$) normal-$\to$ P ($\mathbb{Z}$NegSuc n)) normal-$\to$
  (x : $\mathbb{Z}$) normal-$\to$ P x
$\iota\mathbb{Z}$Pos : (P : $\mathbb{Z}$ normal-$\to$ Set) normal-$\to$
  (posP : (n : $\mathbb{N})$ normal-$\to$
    P ($\mathbb{Z}$Pos n)) normal-$\to$
  (negSucP : (n : $\mathbb{N})$ normal-$\to$
    P ($\mathbb{Z}$NegSuc n)) normal-$\to$
  (n : $\mathbb{N}$) normal-$\to$
  (Q : P ($\mathbb{Z}$Pos normal-$\to$ Set) normal-$\to$
  Q ($\mathbb{Z}$ P posP negSucP
      ($\mathbb{Z}$Pos n)) normal-$\to$
  Q (posP n)
$\iota\mathbb{Z}$NegSuc : (P : $\mathbb{Z}$ normal-$\to$ Set) normal-$\to$
  (posP : (n : $\mathbb{N}$) normal-$\to$
    P ($\mathbb{Z}$Pos n)) normal-$\to$
  (negSucP : (n : $\mathbb{N})$ normal-$\to$
    P ($\mathbb{Z}$NegSuc n)) normal-$\to$
  (n : $\mathbb{N}$) normal-$\to$
  (Q : P ($\mathbb{Z}$NegSuc normal-$\to$ Set) normal-$\to$
  Q ($\mathbb{Z}$ P posP negSucP
      ($\mathbb{Z}$NegSuc n)) normal-$\to$
  Q (negSucP n)
depConstrGZPos : $\mathbb{N}$ normal-$\to$ GZ
depConstrGZPos n = [ (n , 0) ]
depConstrGZNegSuc : $\mathbb{N}$ normal-$\to$ GZ
depConstrGZNegSuc n = [ (0 , suc n) ]
depElimGZ : (P : GZ normal-$\to$ Type) normal-$\to$
  ((x : GZ) normal-$\to$ isSet (P x)) normal-$\to$
  ((n : $\mathbb{N})$ normal-$\to$ P (depConstrGZPos n)) normal-$\to$
  ((n : $\mathbb{N})$ normal-$\to$ P (depConstrGZNegSuc n)) normal-$\to$
  (x : GZ) normal-$\to$ P x
$\iota$GZPos : (P : GZ normal-$\to$ Set) normal-$\to$
  ((x : GZ) normal-$\to$ isSet (P x)) normal-$\to$
  (posP : (n : $\mathbb{N})$ normal-$\to$
    P (depConstrGZPos n)) normal-$\to$
  (negSucP : (n : $\mathbb{N})$ normal-$\to$
    P (depConstrGZNegSuc n)) normal-$\to$
  (n : $\mathbb{N}$) normal-$\to$
  (Q : P (depConstrGZPos normal-$\to$ Set) normal-$\to$
  Q (depElimGZ P posP negSucP
      (depConstrGZPos n)) normal-$\to$
  Q (posP n)
$\iota$GZNegSuc : (P : GZ normal-$\to$ Set) normal-$\to$
  ((x : GZ) normal-$\to$ isSet (P x)) normal-$\to$
  (posP : (n : $\mathbb{N}$) normal-$\to$
    P (depConstrGZPos n)) normal-$\to$
  (negSucP : (n : $\mathbb{N})$ normal-$\to$
    P (depConstrGZNegSuc n)) normal-$\to$
  (n : $\mathbb{N}$) normal-$\to$
  (Q : P (depConstrGZNegSuc normal-$\to$ Set) normal-$\to$
  Q (depElimGZ P posP negSucP
      (depConstrGZNegSuc n)) normal-$\to$
  Q (negSucP n)
Figure 15. Configuration for the standard library representation of the integers (left) and the Grothendieck group completion of the integers (right) in Cubical Agda. For clarity, only the types are shown for the eliminators and $\iota$ rules; full proof terms are in the artifact. See Figure 16 for the corresponding reversed $\iota$ rules, denoted with a “${}^{-}$”.
$\iota$$\mathbb{Z}$${}^{-}$ : (P : $\mathbb{Z}$ normal-$\to$ Set) normal-$\to$
  (posP : (n : $\mathbb{N}$) normal-$\to$
    P ($\mathbb{Z}$Pos n)) normal-$\to$
  (negSucP : (n : $\mathbb{N}$) normal-$\to$
    P ($\mathbb{Z}$NegSuc n)) normal-$\to$
  (n : $\mathbb{N}$) normal-$\to$
  (Q : P ($\mathbb{Z}$Pos n) normal-$\to$ Set) normal-$\to$
  Q (posP n) normal-$\to$
  Q ($\mathbb{Z}$ P posP negSucP
      ($\mathbb{Z}$Pos n))
$\iota$$\mathbb{Z}$${}^{-}$ : (P : $\mathbb{Z}$ normal-$\to$ Set) normal-$\to$
  (posP : (n : $\mathbb{N}$) normal-$\to$
    P ($\mathbb{Z}$Pos n)) normal-$\to$
  (negSucP : (n : $\mathbb{N}$) normal-$\to$
    P ($\mathbb{Z}$NegSuc n)) normal-$\to$
  (n : $\mathbb{N}$) normal-$\to$
  (Q : P ($\mathbb{Z}$NegSuc n) normal-$\to$ Set) normal-$\to$
  Q (negSucP n) normal-$\to$
  Q ($\mathbb{Z}$ P posP negSucP
      ($\mathbb{Z}$NegSuc n))
$\iota$${}^{-}$ : (P : GZ normal-$\to$ Set) normal-$\to$   (pset : $\forall$ x normal-$\to$ isSet (P x)) normal-$\to$   (posP : (n : $\mathbb{N}$) normal-$\to$     P (depConstrGZPos n)) normal-$\to$   (negSucP : (n : $\mathbb{N}$) normal-$\to$     P (depConstrGZNegSuc n)) normal-$\to$   (n : $\mathbb{N}$) normal-$\to$   (Q : P (depConstrGZPos n) normal-$\to$ Set) normal-$\to$   Q (posP n) normal-$\to$   Q (depElimGZ P pset posP negSucP       (depConstrGZPos n)) $\iota$${}^{-}$ : (P : GZ normal-$\to$ Set) normal-$\to$   (pset : $\forall$ x normal-$\to$ isSet (P x)) normal-$\to$   (posP : (n : $\mathbb{N}$) normal-$\to$     P (depConstrGZPos n)) normal-$\to$   (negSucP : (n : $\mathbb{N}$) normal-$\to$     P (depConstrGZNegSuc n)) normal-$\to$   (n : $\mathbb{N}$) normal-$\to$   (Q : P (depConstrGZNegSuc n) normal-$\to$ Set) normal-$\to$   Q (negSucP n) normal-$\to$   Q (depElimGZ P pset posP negSucP       (depConstrGZNegSuc n))
Figure 16. The remaining $\iota$ rules for the standard library representation of the integers (left) and the Grothendieck group completion representation of the integers (right) in Cubical Agda. The terms are suffixed by “${}^{-}$” to indicate that they apply an $\iota$-reduction step in reverse.
Definition depConstrZPos (n : nat) : Z := pos n.
Definition depConstrZNegSuc (n : nat) : Z := negsuc n.
Definition depRecZ (C : Type)
  (posP : forall (n : nat), C)
  (negSucP : forall (n : nat), C)
  (z : Z) :
  C
Definition depElimPropZ (P : Z -> Prop)
  (posP : forall (n : nat),
    P (depConstrZPos n))
  (negSucP : forall (n : nat),
    P (depConstrZNegSuc n))
  (z : Z) :
  P z.
Theorem iotaZPos (C : Type)
  (posP : forall (n : nat), C)
  (negSucP : forall (n : nat), C)
  (n : nat) :
  forall (Q : C -> Type),
    (Q (depRecZ C posP negSucP
         (depConstrZPos n))) ->
    Q (posP n).
Theorem iotaZNegSuc (C : Type)
  (posP : forall (n : nat), C)
  (negSucP : forall (n : nat), C)
  (n : nat) :
  forall (Q : C -> Type),
    (Q (depRecZ C posP negSucP
         (depConstrZNegSuc n))) ->
    Q (negSucP n).
Definition depConstrGZPos (n : nat) : GZ := (n, 0).
Definition depConstrGZNegSuc (n : nat) : GZ := (0, S n).
Definition depRecGZ (C : Type)
  (posP : forall (n : nat), C)
  (negSucP : forall (n : nat), C)
  (z : GZ) :
  C.
Theorem depElimPropGZ (P : GZ -> Prop)
  ‘(p : Proper (GZ -> Prop) (eq_GZ ==> iff) P)
  (posP : forall (n : nat),
    P (depConstrGZPos n))
  (negSucP : forall (n : nat),
    P (depConstrGZNegSuc n))
  (z : GZ) :
  P z.
Definition iotaRecGZPos (C : Type)
  (posP : forall (n : nat), C)
  (negSucP : forall (n : nat), C)
  (n : nat) :
  forall (Q : C -> Type),
    (Q (depRecGZ C posP negSucP
         (depConstrGZPos n))) ->
    Q (posP n).
Definition iotaRecGZNegSuc (C : Type)
  (posP : forall (n : nat), C)
  (negSucP : forall (n : nat), C)
  (n : nat) :
  forall (Q : C -> Type),
    (Q (depRecGZ C posP negSucP
         (depConstrGZNegSuc n))) ->
    Q (negSucP n).
Figure 17. Configuration for Cubical Agda’s standard library representation of the integers (left) and the Grothendieck group completion representation of the integers (right) in Coq. For clarity, only the types are shown for the eliminators and $\iota$ rules; full proof terms are in the artifact. See Figure 18 for the corresponding reversed $\iota$ rules.
Theorem iotaZPosRev (C : Type)
  (posP : forall (n : nat), C)
  (negSucP : forall (n : nat), C)
  (n : nat) :
  forall (Q : C -> Type),
  Q (posP n) -> (Q (depRecZ C posP negSucP (depConstrZPos n))).
Theorem iotaZPosRev (C : Type)
  (posP : forall (n : nat), C)
  (negSucP : forall (n : nat), C)
  (n : nat) :
  forall (Q : C -> Type),
  Q (posP n) -> (Q (depRecZ C posP negSucP (depConstrZPos n))).
Definition iotaRecGZPosRev (C : Type)
  (posP : forall (n : nat), C)
  (negSucP : forall (n : nat), C)
  (n : nat) :
  forall (Q : C -> Type),
    Q (posP n) -> (Q (depRecGZ C posP negSucP (depConstrGZPos n))).
Definition iotaRecGZNegSucRev (C : Type)
  (posP : forall (n : nat), C)
  (negSucP : forall (n : nat), C)
  (n : nat) :
  forall (Q : C -> Type),
    Q (negSucP n) -> (Q (depRecGZ C posP negSucP (depConstrGZNegSuc n))).
Figure 18. The remaining $\iota$ rules for Cubical Agda’s standard library representation of the integers (left) and the Grothendieck group completion representation of the integers (right) in Coq.
Appendix B Complete configuration for queue case study
To perform repair between our types of one list queues and two list queues, we need to define the configuration to repair across. In Figures 19 and 20, we can see types of the elements of the configuration defined in Cubical Agda. The left types are the components corresponding to one list queues, and the right types correspond to two list queues. The types are aligned to highlight that the types have the same shape, except for our hypothesis (pset : (q : TLQ) normal-$\to$ isSet (P q)) in depElimTLQ. Notice that there is only one eliminator which can be used with any motive.
In Figures 21 and 22, we see the types of the elements of the configuration defined in Coq. Again, the left types correspond to one list queues, and the right types correspond to two list queues, with the types aligned to highlight that the types have the same shape, except for the hypothesis that:
  ‘(p : Proper (TLQ -> Prop) (eq_queue ==> iff) P)
in depElimProp. Here, there are two eliminators. One eliminates into Type, but does not support dependently typed output, while the other only eliminates into Prop but does support dependently typed output.
depConstrOLQEmpty : OLQ
depConstrOLQInsert : A normal-$\to$ OLQ normal-$\to$ OLQ
depElimOLQ : (P : OLQ normal-$\to$ Type) normal-$\to$
  (P depConstrOLQEmpty) normal-$\to$
  ($\forall$ q a normal-$\to$ (P q) normal-$\to$
    P (depConstrOLQInsert a q)) normal-$\to$
  (x : OLQ) normal-$\to$ P x
$\iota$OLQEmpty : (P : OLQ normal-$\to$ Set) normal-$\to$
  (emptyP : P depConstrOLQEmpty) normal-$\to$
  (insertP :
    (q : OLQ) normal-$\to$ (a : A) normal-$\to$ (P q) normal-$\to$
    P (depConstrOLQInsert a q)) normal-$\to$
  (Q : P depConstrOLQEmpty normal-$\to$ Set) normal-$\to$
  Q (depElimOLQ P emptyP insertP depConstrOLQEmpty) normal-$\to$
  Q emptyP
$\iota$OLQInsert : (P : OLQ normal-$\to$ Set) normal-$\to$
  (emptyP : P depConstrOLQEmpty) normal-$\to$
  (insertP : (q : OLQ) normal-$\to$ (a : A) normal-$\to$
    (P q) normal-$\to$
    P (depConstrOLQInsert a q)) normal-$\to$
  (a : A) normal-$\to$ (q : OLQ) normal-$\to$
  (Q : P (depConstrOLQInsert a q) normal-$\to$
       Set) normal-$\to$
  Q (depElimOLQ P emptyP insertP
      (depConstrOLQInsert a q)) normal-$\to$
  Q (insertP q a (depElimOLQ P emptyP insertP q))
depConstrTLQEmpty : TLQ
depConstrTLQInsert : A normal-$\to$ TLQ normal-$\to$ TLQ
depElimTLQ : (P : TLQ normal-$\to$ Type) normal-$\to$
  ($\forall$ x normal-$\to$ isSet (P x)) normal-$\to$
  (P depConstrTLQEmpty) normal-$\to$
  ($\forall$ q a normal-$\to$ (P q) normal-$\to$
    P (depConstrTLQInsert a q)) normal-$\to$
  (x : TLQ) normal-$\to$ P x
$\iota$TLQEmpty : (P : TLQ normal-$\to$ Set) normal-$\to$
  (pset : (q : TLQ) normal-$\to$ isSet (P q))
  (emptyP : P depConstrTLQEmpty) normal-$\to$
  (insertP : (q : TLQ) normal-$\to$ (a : A) normal-$\to$
    (P q) normal-$\to$
    P (depConstrTLQInsert a q)) normal-$\to$
  (Q : P depConstrTLQEmpty normal-$\to$ Set) normal-$\to$
  Q (depElimTLQ P emptyP insertP depConstrTLQEmpty) normal-$\to$
  Q emptyP
$\iota$TLQInsert : (P : TLQ normal-$\to$ Set) normal-$\to$
  (pset : (q : TLQ) normal-$\to$ isSet (P q))
  (emptyP : P depConstrTLQEmpty) normal-$\to$
  (insertP : (q : TLQ) normal-$\to$ (a : A) normal-$\to$
    (P q) normal-$\to$
    P (depConstrTLQInsert a q)) normal-$\to$
  (a : A) normal-$\to$ (q : TLQ) normal-$\to$
  (Q : P (depConstrTLQInsert a q) normal-$\to$
       Set) normal-$\to$
  Q (depElimTLQ P emptyP insertP
      (depConstrTLQInsert a q)) normal-$\to$
  Q (insertP q a (depElimTLQ P emptyP insertP q))
Figure 19. Configuration for one list queues (left) and two list queues (right) in Cubical Agda. For clarity, only the types are shown; full proof terms are in the artifact. See Figure 20 for the corresponding reversed $\iota$ rules, denoted with a “${}^{-}$”.
$\iota$${}^{-}$ : (P : OLQ normal-$\to$ Set) normal-$\to$
  (emptyP : P depConstrOLQEmpty) normal-$\to$
  (insertP : (q : OLQ) normal-$\to$ (a : A) normal-$\to$
    (P q) normal-$\to$ P (depConstrOLQInsert a q)) normal-$\to$
  (Q : P (depConstrOLQEmpty) normal-$\to$ Set) normal-$\to$
  Q emptyP normal-$\to$
  Q (depElimOLQ P emptyP insertP
       depConstrOLQEmpty)
$\iota$${}^{-}$ : (P : OLQ normal-$\to$ Set) normal-$\to$
  (emptyP : P depConstrOLQEmpty) normal-$\to$
  (insertP : (q : OLQ) normal-$\to$ (a : A) normal-$\to$
    (P q) normal-$\to$ P (depConstrOLQInsert a q)) normal-$\to$
  (a : A) normal-$\to$ (q : OLQ) normal-$\to$
  (Q : P (depConstrOLQInsert a q) normal-$\to$
       Set) normal-$\to$
  Q (insertP q a (depElimOLQ P emptyP
      insertP q)) normal-$\to$
  Q (depElimOLQ P emptyP insertP
      (depConstrOLQInsert a q))
$\iota$${}^{-}$ : (P : TLQ normal-$\to$ Set) normal-$\to$
  (pset : (q : TLQ) normal-$\to$ isSet (P q)) normal-$\to$
  (emptyP : P depConstrTLQEmpty) normal-$\to$
  (insertP : (q : TLQ) normal-$\to$ (a : A) normal-$\to$
    (P q) normal-$\to$ P (depConstrTLQInsert a q)) normal-$\to$
  (Q : P depConstrTLQEmpty normal-$\to$ Set) normal-$\to$
  Q emptyP normal-$\to$
  Q (depElimTLQ P pset emptyP insertP depConstrTLQEmpty)
$\iota$${}^{-}$ : (P : TLQ normal-$\to$ Set) normal-$\to$
  (pset : (q : TLQ) normal-$\to$ isSet (P q)) normal-$\to$
  (emptyP : P depConstrTLQEmpty) normal-$\to$
  (insertP : (q : TLQ) normal-$\to$ (a : A) normal-$\to$
    (P q) normal-$\to$ P (depConstrTLQInsert a q)) normal-$\to$
  (a : A) normal-$\to$ (q : TLQ) normal-$\to$
  (Q : P (depConstrTLQInsert a q) normal-$\to$
       Set) normal-$\to$
  Q (insertP q a (depElimTLQ P pset emptyP
      insertP q)) normal-$\to$
  Q (depElimTLQ P pset emptyP insertP
      (depConstrTLQInsert a q))
Figure 20. The remaining $\iota$ rules for one list queues (left) and two list queues (right) in Cubical Agda. The terms are suffixed by “${}^{-}$” to indicate that they apply an $\iota$-reduction step in reverse.
Definition depConstrOLQEmpty : OLQ := [].
Definition depConstrOLQInsert (a : A)
  (q : OLQ) : OLQ :=
  a :: q.
Definition depRecOLQ (C : Type)
  (pEmpty : C)
  (pInsert : forall (a : A) (q : OLQ),
    C -> C) :
  (forall (x : OLQ), C).
Definition depElimPropOLQ (P : OLQ -> Prop)
  (pEmpty : P depConstrOLQEmpty)
  (pInsert : forall (a : A) (q : OLQ),
    P q -> P (depConstrOLQInsert a q)) :
  (forall (x : OLQ), P x).
Theorem iotaRecOLQEmpty (C : Type)
  (pEmpty : C)
  (pInsert : forall (a : A) (q : OLQ),
    C -> C) :
  forall (Q : C -> Type),
    (Q (depRecOLQ C
      pEmpty
      pInsert
      depConstrOLQEmpty)) ->
    (Q pEmpty).
Theorem iotaRecOLQInsert (C : Type)
  (pEmpty : C)
  (pInsert : forall (a : A) (q : OLQ),
    C -> C)
  (a : A)
  (q : OLQ) :
forall (a : A) (q : OLQ) (Q : C -> Type),
  Q (depRecOLQ C pEmpty pInsert
      (depConstrOLQInsert a q)) ->
  Q (pInsert a q (depRecOLQ C pEmpty pInsert q)).
Definition depConstrTLQEmpty : TLQ := ([],[]).
Definition depConstrTLQInsert (a : A)
  (q : TLQ) : TLQ :=
  match q with
  | (l1, l2) => (a :: l1, l2)
  end.
Definition depRecTLQ (C : Type)
  (pEmpty : C)
  (pInsert : forall (a : A) (q : TLQ), C -> C) :
  (forall (x : TLQ), C).
Theorem depElimPropTLQ (P : TLQ -> Prop)
  ‘(p : Proper (TLQ -> Prop) (eq_queue ==> iff) P)
  (pEmpty : P depConstrTLQEmpty)
  (pInsert : forall (a : A) (q : TLQ),
    P q -> P (depConstrTLQInsert a q)) :
  (forall (x : TLQ), P x).
Theorem iotaRecTLQEmpty (C : Type)
  (pEmpty : C)
  (pInsert : forall (a : A) (q : TLQ),
    C -> C) :
forall (Q : C -> Type),
    (Q (depRecTLQ C
      pEmpty
      pInsert
      depConstrTLQEmpty)) ->
    (Q pEmpty).
Theorem iotaRecTLQInsert (C : Type)
  (pEmpty : C)
  (pInsert : forall (a : A) (q : TLQ),
    C -> C)
  (a : A)
  (q : TLQ) :
forall (a : A) (q : TLQ) (Q : C -> Type),
  Q (depRecTLQ C pEmpty pInsert
      (depConstrTLQInsert a q)) ->
  Q (pInsert a q (depRecTLQ C pEmpty
       pInsert q)).
Figure 21. Configuration for one list queues (left) and two list queues (right) in Coq. For clarity, only the types are shown for the eliminators and $\iota$ rules; full proof terms are in the artifact. See Figure 22 for the corresponding reversed $\iota$ rules.
Theorem iotaRecOLQEmptyRev (C : Type)
  (pEmpty : C)
  (pInsert : forall (a : A) (q : OLQ),
    C -> C) :
  forall (Q : C -> Type),
    (Q pEmpty) ->
    (Q (depRecOLQ C pEmpty pInsert depConstrOLQEmpty)).
Theorem iotaRecOLQInsertRev (C : Type)
  (pEmpty : C)
  (pInsert : forall (a : A) (q : OLQ),
    C -> C) (a : A) (q : OLQ) :
forall (a : A) (q : OLQ) (Q : C -> Type),
  Q (pInsert a q (depRecOLQ
      C pEmpty pInsert q)) ->
  Q (depRecOLQ
      C pEmpty pInsert
      (depConstrOLQInsert a q)).
Theorem iotaRecTLQEmptyRev (C : Type)
  (pEmpty : C)
  (pInsert : forall (a : A) (q : TLQ),
    C -> C) :
  forall (Q : C -> Type),
    (Q pEmpty) ->
    (Q (depRecTLQ C pEmpty pInsert depConstrTLQEmpty)).
Theorem iotaRecTLQInsertRev (C : Type)
  (pEmpty : C)
  (pInsert : forall (a : A) (q : TLQ),
    C -> C) (a : A) (q : TLQ) :
forall (a : A) (q : TLQ) (Q : C -> Type),
  Q (pInsert a q (depRecTLQ
    C pEmpty pInsert q)) ->
  Q (depRecTLQ
      C pEmpty pInsert
      (depConstrTLQInsert a q)).
Figure 22. The remaining $\iota$ rules for one list queues (left) and two list queues (right) in Coq.
Appendix C Original and Repaired Dequeue Implementations
Here, we provide the full terms of our dequeue functions, both before and after repair. In Figure 23, we see the definitions in Cubical Agda. The original dequeue implementation is on the left, and the repaired version is on the right. We can see that the repaired implementation is the result of the transformation we define in this paper.
In Figure 24 we see the definitions of dequeue in Coq. The repaired terms on the right were were obtained automatically by running our repair prototype on the original definitions found on the left. Again, we can see that these terms are the result of the transformation we define.
dequeueOLQ : OLQ normal-$\to$ Maybe (OLQ $\times$ A)
dequeueOLQ = depElimOLQ
  ($\lambda$ normal-$\to$ Maybe (OLQ $\times$ A))
  nothing
  recCase where
  recCase: (q : OLQ) (outer : A) normal-$\to$
    Maybe (OLQ $\times$ A) normal-$\to$
    Maybe (OLQ $\times$ A)
  recCase q outer =
    Cubical.Data.Maybe.rec
      (just (depConstrOLQEmpty , outer))
      ($\lambda$ p normal-$\to$ just (depConstrOLQInsert
[…]
dequeueTLQ : TLQ normal-$\to$ Maybe (TLQ $\times$ A)
dequeueTLQ = depElimTLQ
  ($\lambda$ normal-$\to$ Maybe (TLQ $\times$ A))
  ($\lambda$ normal-$\to$ isSetDeqReturnType)
  nothing
  recCase where
  recCase: (q : TLQ) (outer : A) normal-$\to$
    Maybe (TLQ $\times$ A) normal-$\to$
    Maybe (TLQ $\times$ A)
  recCase q outer =
    Cubical.Data.Maybe.rec
      (just (depConstrTLQEmpty , outer))
      ($\lambda$ p normal-$\to$ just (depConstrTLQInsert
[…]
Figure 23. The dequeue function on queues in Cubical Agda. On the left is the original version on one list queues, and on the right is the repaired version on two list queues.
Definition dequeueHelpOLQ (outer : A)
  (q : OLQ) (m : option (OLQ * A)) :
  option (OLQ * A) :=
  @option_rect
    (OLQ * A)
    (fun _ => option (OLQ * A))
    (fun (p : (OLQ * A)) =>
      Some (depConstrOLQInsert
        outer (fst p), (snd p)))
    (Some (depConstrOLQEmpty, outer))
    m.
Definition dequeueOLQ :
  OLQ -> option (OLQ * A) :=
  depRecOLQ (option (OLQ * A)) None dequeueHelpOLQ.
Definition dequeueHelpTLQ : A -> TLQ ->
  option (TLQ * A) -> option (TLQ * A) :=
  fun (outer : A) (_ : TLQ)
    (m : option (TLQ * A)) =>
    option_rect
      (fun _ : option (TLQ * A) =>
        option (TLQ * A))
      (fun p : TLQ * A =>
        Some (depConstrTLQInsert outer (fst p), snd p))
      (Some (depConstrTLQEmpty, outer))
      m
Definition dequeueTLQ :
  TLQ -> option (TLQ * A) :=
  depRecTLQ (option (TLQ * A)) None dequeueHelpTLQ
Figure 24. The dequeue function on queues in Coq. On the left is the original version on one list queues, and on the right is the repaired version on two list queues.